OpenStudy (anonymous):

cuberoot(x+2)=cuberoot(x-2)+2

OpenStudy (anonymous):

do you know?

OpenStudy (anonymous):

yes i am trying on paper to see if i can solve it......

OpenStudy (anonymous):

sorry friend, i am not sure about what i am arriving at, so i guess someone else might have to help you

OpenStudy (anonymous):

i tried it , getting 0=x^2-2x+8, and thats not factorable, quite confused if im doing it right...

OpenStudy (anonymous):

now im getting complex roots...

OpenStudy (anonymous):

actually I live in India Right now it is 2 AM in the morning and the brain is tired as I have been at it for many hours....

OpenStudy (anonymous):

cool, what did you get though?

OpenStudy (anonymous):

I now have +-2i

OpenStudy (anonymous):

is anyone going to verify this?

OpenStudy (nikvist):

\[x_1=\frac{10}{9}\sqrt{3}\,\,,\,\,x_2=-\frac{10}{9}\sqrt{3}\]